# Can Authorship Attribution Models Distinguish Speakers in Speech Transcripts?

### v2

@@ Abstract
Authorship verification is the task of determining if two distinct writing samples share the same author and is typically concerned with the attribution of written text. In this paper, we explore the attribution of transcribed speech, which poses novel challenges. The main challenge is that many stylistic features, such as punctuation and capitalization, are not informative in this setting. On the other hand, transcribed speech exhibits other patterns, such as filler words and backchannels (e.g., um, uh-huh), which may be characteristic of different speakers. We propose a new benchmark for speaker attribution focused on conversational speech transcripts. To limit spurious associations of speakers with topic, we employ both conversation prompts and speakers participating in the same conversation to construct verification trials of varying difficulties. We establish the state of the art on this new benchmark by comparing a suite of neural and non-neural baselines, finding that although written text attribution models achieve surprisingly good performance in certain settings, they perform markedly worse as conversational topic information is increasingly controlled. We present analyses of the impact of transcription style on performance as well as the ability of fine-tuning on speech transcripts to improve performance.111Our benchmark is available at github.com/
@@ 1 Introduction
There are a number of motivations for exploring the use of textual authorship attribution methods on transcribed speech. In challenging acoustic settings, such as with degraded audio or a disguised voice, speech content and syntax may be the only reliable signal available. In other cases, the original audio may no longer be available, leaving only a textual version of the speech. This is common with transcripts of interviews, court proceedings, and in commercial settings where text is preferred for archiving. Developing reliable identification of a speaker based on transcripts of their speech may also expose potential blind spots in current speaker anonymization approaches, which alter the speech signal but leave the speech content intact (Sisman et al., 2021; Fang et al., 2019).
@@
Despite the potential utility of distinguishing speakers via transcribed speech, there are a number of challenges. One that applies to both transcript-based and text-based attribution is the need for substantial writing samples to characterize style, which, in the speech domain, means a large number of utterances. Also, since transcribed speech is both a different domain and a different modality, text-based attribution models have more of a gulf to bridge than in a standard cross-domain task. For example, while written text contains features such as punctuation, spelling, and use of emoticons that can vary systematically across individuals, transcribed speech may lack these particular stylistic cues. Instead, speech contains other potentially identifying features, such as filler words (e.g. um), backchannels (e.g. uh-huh), and other discourse markers (e.g. well, I mean) (Duncan, 1974; Sacks, 1992), and, more rarely, indications of pause length or intonation. Finally, working with transcribed speech, much like working with translated language, introduces noise into the system via the transcription or translation process.
The present work contributes a benchmark for text-based authorship attribution models on conversational speech transcripts. Using the task of verification—i.e. determining if two transcripts have the same speaker or different speakers—we perform a systematic comparison of existing methods and provide a detailed analysis of the results. This work focuses on the following research questions. First, despite the difference in modality, do textual authorship models transfer to speech transcripts? Once we have established this benchmark, we probe other aspects of the task to help determine the limits of authorship models applied to transcripts. Since speech is transcribed based on a transcription style, what is the impact of transcription style on attribution performance? As alluded to above, many state-of-the-art authorship models focus on features such as capitalization that can be erased or standardized with transcription. Additionally, many models rely on topic as a clue for attribution, so to what extent does controlling for topic make the task harder? Next, does fine-tuning on speech transcripts significantly improve performance, and then, does further pre-training on speech transcripts improve it more? Finally, since in many forensic settings there are often limited data, how many utterances are required to achieve a given level of verification performance? By addressing these questions we provide a proof of concept for the viability of applying authorship models to speech transcripts.
@@ 2 Related Work
Early work in speaker identification for text used special-purpose models for novels that match utterances to the character who ‘said’ them (He et al., 2013). The first work we are aware of to perform attribution on speech transcripts is a study that looks at word frequency to distinguish 250 speakers of Dutch (Scheijen, 2020). In contrast, we consider a larger set of speakers and compare a range of different methods, both neural and non-neural. A contemporaneous related task is the PAN 2023 competition, which looked at cross-discourse type authorship verification between essays, emails, interviews, and speech transcripts (Stamatatos et al., 2023). However, since topic is not controlled across the data, it is unclear which features are driving the models’ performance. In contrast, we more carefully control conversation topic in the design of our benchmark since this is a confounding variable in evaluating the ability of attribution methods to capture distinct speaker styles.
Recently, Tripto et al. (2023) compare a number of statistical and neural authorship models on human spoken texts and large language models prompted to emulate spoken texts, finding that even simple -gram-based authorship models can perform well on speech transcripts. However, we present contradictory results in this work, finding that most text-based authorship models have almost no predictive power once topic is controlled for. We conduct further experiments that tease apart which factors influence performance, such as transcription style and number of utterances in the transcript.
@@ 3 A Speaker Attribution Benchmark
To compare performance across the range of intended conditions, we focus on one dataset that fit all our requirements, namely large enough in number of speakers, number of utterances per speaker, and number of conversations to allow fine-tuning. We also focus on gold standard transcriptions since our aim is to establish an initial benchmark of performance. To be able to adjust the difficulty level of the verification task to obtain a range of performance for the models, it was crucial that speakers discuss fixed topics, which would allow matching transcripts not only based on speaker, but also on the content of what is discussed. Finally, a conversational dataset aligns with many likely use cases of speaker attribution, especially in the forensic setting.
We chose the Fisher English Training Speech Transcripts corpus (Cieri et al., 2004), a collection of transcribed phone calls, due to its accessibility, size, manual annotation of errors, conversational topic assignments, and gender balance amongst speakers. The corpus consists of 11,699 transcribed phone calls totalling 1,960 hours. Participants on calls generally did not know each other and had an assigned discussion topic that was randomly selected from a list. In general, participants stayed on topic throughout the duration of the call (Cieri et al., 2004), which we confirmed through a manual check of a random sample of the transcripts. The calls lasted 10 minutes and speakers often participated in multiple calls.
The Fisher corpus contains two transcript ‘encodings’, or annotation styles. One was transcribed using WordWave quick transcription with error corrections and post-processing by BBN Technologies. This ‘BBN’ encoding includes prescriptive punctuation and capitalization according to an existing WordWave style guide (Kimball et al., n.d.). The Linguistic Data Consortium (LDC) provided the second encoding, with automatic segmentation of the audio data and manual transcription of the words, including a basic spell check (Cieri et al., 2004). This transcription did not include punctuation (other than apostrophes and hyphens), put text in all lowercase, and often grouped together text by the same speaker despite interjected backchannels. Comparing performance on each encoding can help elucidate the extent to which the models capture ‘deeper’ authorship features rather than surface-level low-hanging fruit. Figure 1 presents an example of each. Both encodings include non-speech sounds, such as laughing and undistinguished noise, in square brackets. The LDC encoding employs double parentheses to identify unclear productions that were guessed at by the transcriber.
@@
We select the same transcripts from both encodings to test the impact of annotation on a model’s attribution performance, according to the following procedure. First, we use a 50/25/25 training/validation/test split with no overlap in speakers between training and evaluation splits, making the task more challenging. For the speakers in each set, verification trials are formed by matching two transcripts from either the same speaker (‘positive’) or different speakers (‘negative’). Transcripts have ~1400 tokens across an average of 100 utterances. In Section 4.6 we vary the number of utterances used from each transcript; the results revealed that the appearance of names during introductions at the beginning of the call significantly helped model performance; as a result, we remove the first five utterances of each transcript.222Many introductions concluded within the first two utterances per speaker but we removed five to be conservative. We also ran a simple check for (re)introductions later in the transcript but found them to be rare.
@@ 4 Experiments
Models We test and compare the performance of four main models. The first is Sentence-BERT (SBERT)333huggingface.co/sentence-transformers/all-MiniLM-L12-v2, a variant of the pretrained BERT network that creates semantically-related sentence embeddings (Reimers and Gurevych, 2019). As a complement to the content-focused SBERT, we test Content-Independent Style Representations (CISR)444huggingface.co/AnnaWegmann/Style-Embedding, which aims to capture writing style rather than content by controlling the topic of verification trials at training time (Wegmann et al., 2022). The third model is an instance of Learning Universal Authorship Representations (LUAR)555huggingface.co/rrivera1849/LUAR-MUD, which does well with zero-shot transfer between Reddit, Amazon, and fanfiction stories (Rivera-Soto et al., 2021), capturing stylistic features of an author’s writing with less content-sensitivity (Wang et al., 2023). In summary, we use two models tailored to capture primarily content (SBERT) and primarily style (CISR), in addition to a model that combines both and does domain transfer well (LUAR). We specifically use a LUAR model trained on a large Reddit dataset consisting of comments by one million authors (Khan et al., 2021), which we hypothesize is more similar to the conversational speech transcripts than other more formal domains.
The final model we consider is the recurrent neural network model AdHominem666github.com/boenninghoff/AdHominem (Boenninghoff et al., 2019), which performed well on speech transcripts in recent work (Tripto et al., 2023). This model uses a hierarchical architecture to aggregate character, word, and sentence features from each document. For reference, we also include two baselines: TF-IDF, a weighted measure of word overlap, and PAN’s authorship verification baseline of TF-IDF-weighted character 4-grams (PANgrams)777github.com/pan-webis-de/pan-code/tree/master/clef23/authorship-verification (Stamatatos et al., 2023).
@@ 4.1 Experiment 1: Baselines
TF-IDF | 0.537 | 0.528 | 0.489 | 0.534 | 0.532 | 0.488 | 0.522 | 0.513 | 0.489 |
PANgrams | 0.796 | 0.699 | 0.547 | 0.805 | 0.710 | 0.572 | 0.790 | 0.691 | 0.564 |
AdHominem | 0.565 | 0.553 | 0.492 | 0.600 | 0.574 | 0.536 | 0.588 | 0.589 | 0.545 |
SBERT | 0.646 | 0.483 | 0.322 | 0.653 | 0.456 | 0.283 | 0.621 | 0.514 | 0.174 |
CISR | 0.612 | 0.578 | 0.534 | 0.680 | 0.664 | 0.646 | 0.622 | 0.532 | 0.493 |
LUAR | 0.714 | 0.633 | 0.472 | 0.803 | 0.711 | 0.547 | 0.837 | 0.722 | 0.543 |
To address whether text-based authorship models transfer to transcribed speech, we evaluate the baselines (TF-IDF, PANgrams) and main models (AdHominem, SBERT, CISR, LUAR) out-of-the-box on the ‘base’ difficulty level verification trials. Recall that each transcript originally had ~1400 tokens and ~100 utterances on average, but we removed the first five utterances per speaker as we confirmed they contain name and topic introductions in most transcripts.
For the TF-IDF baseline, the vectorizer was fit to the Reuters-21578 corpus, which contains 10,788 news documents and totals 1.3 million words (Lewis, 1997),888We also tried fitting the vectorizer to the training set transcripts but found similar results so fit to Reuters to focus on out-of-the-box performance of written text models. the document-term matrix obtained for each transcript in the test set trials, and the cosine similarity calculated between each matrix in each trial. For the out-of-the-box PAN baseline, PANgrams was trained on the most recent openly available PAN authorship verification dataset, fanfiction stories (Bevendorff et al., 2020), and evaluated on the test set trials using cosine similarity. For the main models, we first obtain an embedding for each transcript in the trial,999While LUAR is hierarchical, accepting multiple utterances as independent inputs, SBERT and CISR encode representations on the individual sentence (or document) level. To accommodate multiple inputs with SBERT and CISR, we embed each utterance independently and obtain the final embedding using the coordinate-wise mean of all utterance vectors. We found that this typically worked better than concatenating the text of all utterances and producing a single embedding. then calculate the cosine similarity between each set of two embeddings. AdHominem is trained using Euclidean distance, so we negate it to compute speaker similarity between pairs of transcripts. We evaluate performance by bootstrapping the area under the receiver operating characteristic curve (AUC) score over 1000 samples of the data.
@@
AUC score test set results are in Table 2. Focusing on the leftmost column (BBN ‘text-like’ encoding, ‘base’ difficulty), there is a clear ranking in performance: PANgrams achieves the highest performance followed by LUAR, SBERT, and CISR, with all scores well above chance (AUC = ). AdHominem and TF-IDF perform worst but still above chance. These initial results—considering only the ‘base’ setting for now—suggest at least some transfer from the text domain to the speech domain. However, we revisit this idea in Section 4.3 when discussing results on topic-controlled datasets.

### v3

@@ Abstract
Authorship verification is the task of determining if two distinct writing samples share the same author and is typically concerned with the attribution of written text. In this paper, we explore the attribution of transcribed speech, which poses novel challenges. The main challenge is that many stylistic features, such as punctuation and capitalization, are not informative in this setting. On the other hand, transcribed speech exhibits other patterns, such as filler words and backchannels (e.g., um, uh-huh), which may be characteristic of different speakers. We propose a new benchmark for speaker attribution focused on human-transcribed conversational speech transcripts. To limit spurious associations of speakers with topic, we employ both conversation prompts and speakers participating in the same conversation to construct verification trials of varying difficulties. We establish the state of the art on this new benchmark by comparing a suite of neural and non-neural baselines, finding that although written text attribution models achieve surprisingly good performance in certain settings, they perform markedly worse as conversational topic is increasingly controlled. We present analyses of the impact of transcription style on performance as well as the ability of fine-tuning on speech transcripts to improve performance.111Our benchmark is available at github.com/
@@ 1 Introduction
There are a number of motivations for exploring the use of textual authorship attribution methods on transcribed speech. In challenging acoustic settings, such as with degraded audio or a disguised voice, speech content and syntax may be the only reliable signal available. In other cases, the original audio may no longer be available, leaving only a textual version of the speech. This is common with transcripts of interviews, court proceedings, and in commercial settings where text is preferred for archiving. Developing reliable identification of a speaker based on transcripts of their speech may also expose potential blind spots in current speaker anonymization approaches, which alter the speech signal but leave the speech content intact (Fang et al., 2019; Sisman et al., 2021).
@@
Despite the potential utility of distinguishing speakers via transcribed speech, there are a number of challenges. One that applies to both transcript-based and text-based attribution is the need for substantial writing samples to characterize style, which, in the speech domain, means a large number of utterances. Also, since transcribed speech is both a different domain and a different modality,222Although transcribed speech is in text form, it still originates from the modality of speech and is thus sufficiently different from other text domains. text-based attribution models have more of a gulf to bridge than in a standard cross-domain task. For example, while written text contains features such as punctuation, spelling, and use of emoticons that can vary systematically across individuals, transcribed speech may lack these particular stylistic cues. Instead, speech contains other potentially identifying features, such as filler words (e.g. um), backchannels (e.g. uh-huh), and other discourse markers (e.g. well, I mean) (Duncan, 1974; Sacks, 1992), and, more rarely, indications of pause length or intonation. Finally, working with transcribed speech, much like working with translated language, introduces noise into the system via the transcription or translation process.
The present work contributes a benchmark for text-based authorship attribution models on human-transcribed conversational speech transcripts. Using the task of verification—i.e. determining if two transcripts have the same speaker or different speakers—we perform a systematic comparison of existing methods and provide a detailed analysis of the results. This work focuses on the following research questions. First, despite the difference in modality, do textual authorship models transfer to speech transcripts? Once we have established this benchmark, we probe other aspects of the task to help determine the limits of authorship models applied to transcripts. Since speech is transcribed based on a transcription style, what is the impact of transcription style on attribution performance? As alluded to above, many state-of-the-art authorship models focus on features such as capitalization that can be erased or standardized with transcription. Additionally, many models rely on topic as a clue for attribution, so to what extent does controlling for topic make the task harder? Next, does fine-tuning on speech transcripts significantly improve performance, and then, does further pre-training on speech transcripts improve it more? Finally, since in many forensic settings there are often limited data, how many utterances are required to achieve a given level of verification performance? By addressing these questions we provide a proof of concept for the viability of applying authorship models to speech transcripts.
@@ 2 Related Work
Early work in speaker identification for text used special-purpose models for novels that match utterances to the character who ‘said’ them (He et al., 2013). The first work we are aware of to perform attribution on speech transcripts is a study that looks at word frequency to distinguish 250 speakers of Dutch (Scheijen, 2020). In contrast, we consider a larger set of speakers and compare a range of different methods, both neural and non-neural. A contemporaneous related task is the PAN 2023 competition, which looked at cross-discourse type authorship verification between essays, emails, interviews, and speech transcripts (Stamatatos et al., 2023). They perform a weaker version of topic control, though, replacing named entities with generic tags.
Recently, Tripto et al. (2023) compare a number of statistical and neural authorship models on human spoken texts and large language models prompted to emulate spoken texts, finding that even simple -gram-based authorship models can perform well on speech transcripts. However, we present contradictory results in this work, finding that most text-based authorship models have almost no predictive power once topic is controlled. We conduct further experiments that tease apart which factors influence performance, such as transcription style and number of utterances in the transcript.
The effect of conversational topic (and text genre) on text-based authorship attribution performance has been studied particularly in (forensic) stylometry to address cases of a domain mismatch between the texts of unknown authorship that are under investigation (test data) and the available comparison texts of known authorship (training data) (Stamatatos, 2018). For instance, an anonymous social media post might have to be compared to the news articles and blog posts of potential authors. These stylometric studies often focus on small amounts of data and/or few candidate authors, such as manually elicited data (Baayen et al., 2002; Goldstein-Stewart et al., 2009) or select literary authors (Kestemont et al., 2012).
Authorship studies on larger amounts of data across many authors use either no topic control or approximate topic control through domain labels, such as categories of Amazon product reviews (Boenninghoff et al., 2019; Zhu and Jurgens, 2021) and subreddits of Reddit (Wegmann et al., 2022; Zhu and Jurgens, 2021). Although it is unclear how representative domain labels are of various topics, studies that implement some version of topic control generally find that performance decreases as topic control increases (Wegmann et al., 2022).
@@ 3 A Speaker Attribution Benchmark
To compare performance across the range of intended conditions, we focus on one dataset that fit all our requirements, namely large enough in number of speakers, number of utterances per speaker, and number of conversations to allow fine-tuning. We also focus on gold standard transcriptions since our aim is to establish an initial benchmark of performance under ideal conditions that can be used as a reference point for future experiments on noisier data. In fact, experimenting on the noisier output of automatic transcribers is out-of-scope for this paper but is an important next step in future work. To be able to adjust the difficulty level of the verification task to obtain a range of performance for the models, it was crucial that speakers discuss fixed topics, which would allow matching transcripts not only based on speaker, but also on the content of what is discussed. Finally, a conversational dataset aligns with many likely use cases of speaker attribution, especially in the forensic setting.
We chose the Fisher English Training Speech Transcripts corpus (Cieri et al., 2004), a collection of human-transcribed phone calls, due to its accessibility, size, manual error correction, conversational topic assignments, and gender balance amongst speakers. The corpus consists of 11,699 transcribed phone calls totalling 1,960 hours. Participants on calls generally did not know each other and had an assigned discussion topic that was randomly selected from a list. In general, participants stayed on topic throughout the duration of the call (Cieri et al., 2004), which we confirmed through a manual check of a random sample of the transcripts. The calls lasted 10 minutes and speakers often participated in multiple calls.
The Fisher corpus contains two transcript ‘encodings’, or annotation styles. One was manually transcribed using WordWave quick transcription with error corrections and post-processing by BBN Technologies. This ‘BBN’ encoding includes prescriptive punctuation and capitalization according to an existing WordWave style guide (Kimball et al., n.d.). The Linguistic Data Consortium (LDC) provided the second encoding, with automatic segmentation of the audio data and manual transcription of the words, including a basic spell check (Cieri et al., 2004). This transcription did not include punctuation (other than apostrophes and hyphens), put text in all lowercase, and often grouped together text by the same speaker despite interjected backchannels. Comparing performance on each encoding can help elucidate the extent to which the models capture ‘deeper’ authorship features rather than surface-level low-hanging fruit. Figure 1 presents an example of each. Both encodings include non-speech sounds, such as laughing and undistinguished noise, in square brackets. The LDC encoding employs double parentheses for unclear productions that were guessed by the transcriber.
@@
L:
Hi. [LAUGH] So, do you have pets?
R:
Ah, no.
L:
Oh. I ha- –
R:
Do you?
L:
Yeah. I do. I have three dogs [LAUGH] –
R:
Oh, okay.
L:
– and I have a bunch of fish. I have –
R:
Oh.
L:
Yeah. I have – I have a black lab; he’s eighty pounds, big guy. And then I have two little dogs, like terrier mixes [LAUGH].
@@
A:
hi [laughter] so do you have pets
B:
(( ah no ))
A:
oh
A:
i ha- yeah i do i have three dogs [laughter]
B:
(( do you ))
B:
oh okay
A:
and i have a bunch of fish i have yeah i have i have a black lab he’s eighty pounds big guy and then i have two little dogs like terrier mixes
B:
(( oh ))
We select the same transcripts from both encodings to test the impact of annotation on a model’s attribution performance, according to the following procedure. First, we use a 50/25/25 training/validation/test split with no overlap in speakers between training and evaluation splits, making the task more challenging. For the speakers in each set, verification trials are formed by matching two transcripts from either the same speaker (‘positive’) or different speakers (‘negative’). Transcripts have ~1400 tokens across an average of 100 utterances. In Section 4.6 we vary the number of utterances used from each transcript; the results revealed that the appearance of names during introductions at the beginning of the call significantly helped model performance; as a result, we remove the first five utterances of each transcript for all experiments.333Many introductions concluded within the first two utterances per speaker but we removed five to be conservative. We also ran a simple check for (re)introductions later in the transcript but found them to be rare.
@@ 4 Experiments
Models We test and compare the performance of four main models. The first is Sentence-BERT (SBERT),444huggingface.co/sentence-transformers/all-MiniLM-L12-v2 a variant of the pretrained BERT network that creates semantically-related sentence embeddings (Reimers and Gurevych, 2019). As a complement to the content-focused SBERT, we test Content-Independent Style Representations (CISR),555huggingface.co/AnnaWegmann/Style-Embedding which aims to capture writing style rather than content by controlling the topic of verification trials at training time (Wegmann et al., 2022). The third model is an instance of Learning Universal Authorship Representations (LUAR),666huggingface.co/rrivera1849/LUAR-MUD which does well with zero-shot transfer between Reddit, Amazon, and fanfiction stories (Rivera-Soto et al., 2021), capturing stylistic features of an author’s writing with less content-sensitivity (Wang et al., 2023). In summary, we use two models tailored to capture primarily content (SBERT) and primarily style (CISR), in addition to a model that combines both and does domain transfer well (LUAR). We specifically use a LUAR model trained on a large Reddit dataset consisting of comments by one million authors (Khan et al., 2021), which we hypothesize is more similar to the conversational speech transcripts than other more formal domains.
The final model we consider is the recurrent neural network model AdHominem777github.com/boenninghoff/AdHominem (Boenninghoff et al., 2019), which performed well on speech transcripts in recent work (Tripto et al., 2023). This model uses a hierarchical architecture to aggregate character, word, and sentence features from each document. For AdHominem, we converted the transcript trials into the appropriate format and then saved a model checkpoint, whose weights were restored for extracting features from the utterances in each trial.
For reference, we also include two baselines: TF-IDF, a weighted measure of word overlap, and PAN’s authorship verification baseline of TF-IDF-weighted character 4-grams (PANgrams)888github.com/pan-webis-de/pan-code/tree/master/clef23/authorship-verification (Stamatatos et al., 2023).999We did not include the best performers (all SBERT-based models) on the contemporaneous PAN 2023 authorship verification competition, whose task involved written data and transcribed spoken data, because neither the models nor the training data are publicly available.
@@ 4.1 Experiment 1: Baselines
| TF-IDFo | 0.537 | 0.528 | 0.489 | 0.534 | 0.532 | 0.488 | 0.522 | 0.513 | 0.489 |
| PANgramso | 0.796 | 0.699 | 0.547 | 0.805 | 0.710 | 0.572 | 0.790 | 0.691 | 0.564 |
| AdHominemo | 0.565 | 0.553 | 0.492 | 0.600 | 0.574 | 0.536 | 0.588 | 0.589 | 0.545 |
| SBERTo | 0.646 | 0.483 | 0.322 | 0.653 | 0.456 | 0.283 | 0.621 | 0.514 | 0.174 |
| CISRo | 0.612 | 0.578 | 0.534 | 0.680 | 0.664 | 0.646 | 0.622 | 0.532 | 0.493 |
| LUARo | 0.714 | 0.633 | 0.472 | 0.803 | 0.711 | 0.547 | 0.837 | 0.722 | 0.543 |
To address whether text-based authorship models transfer to transcribed speech, we evaluate the baselines (TF-IDFo, PANgramso) and main models (AdHominemo, SBERTo, CISRo, LUARo) out-of-the-box on the ‘base’ difficulty level verification trials. Recall that each transcript originally had ~1400 tokens and ~100 utterances on average, but we removed the first five utterances per speaker as we confirmed they contain name and topic introductions in most transcripts.
For the TF-IDF baseline, the vectorizer was fit to the Reuters-21578 corpus, which contains 10,788 news documents and totals 1.3 million words (Lewis, 1997),101010We also tried fitting the vectorizer to the training set transcripts but found similar results so fit to Reuters to focus on out-of-the-box performance of written text models. the document-term matrix obtained for each transcript in the test set trials, and the cosine similarity calculated between each matrix in each trial. For the out-of-the-box PAN baseline, PANgramso was trained on the most recent openly available PAN authorship verification dataset, fanfiction stories (Bevendorff et al., 2020), and evaluated on the test set trials using cosine similarity.
For the main models, we first obtain an embedding for each transcript in the trial,111111While LUAR is hierarchical, accepting multiple utterances as independent inputs, SBERT and CISR encode representations on the individual sentence (or document) level. To accommodate multiple inputs with SBERT and CISR, we embed each utterance independently and obtain the final embedding using the coordinate-wise mean of all utterance vectors. We found that this typically worked better than concatenating the text of all utterances and producing a single embedding. then calculate the cosine similarity between each set of two embeddings. AdHominem is trained using Euclidean distance, so we negate it to compute speaker similarity between pairs of transcripts.
We evaluate the performance of all models by bootstrapping the area under the receiver operating characteristic curve (AUC) score with 1000 resamples of the test data. For testing statistical significance between the first and second best performers, as indicated in all tables in the paper, we used a paired t-test over 1000 resamples to test the null hypothesis that the AUC (or EER) scores produced by two models are the same.121212We also ran a non-parametric test, the Wilcoxon signed-rank test, but since the results were similar, we report the (more conservative) results from the more powerful parametric paired t-test.
@@
AUC score test set results are in Table 2 and EER results are in Table 8 in Appendix A. Focusing on the leftmost column (BBN ‘text-like’ encoding, ‘base’ difficulty), there is a clear ranking in performance: PANgramso achieves the highest performance followed by LUARo, SBERTo, and CISRo, with all scores well above chance (AUC = ). AdHominemo and TF-IDFo perform worst but still above chance. These initial results—considering only the ‘base’ setting for now—suggest at least some transfer from the text domain to the speech domain. However, we revisit this idea in Section 4.3 when discussing results on topic-controlled datasets.